Incorrect o.oo points out of 4.00 P Flag question Find the area under the standard normal curve over the interval specified below. To the left of z1 Select one a. 0.023 b. 0.999 C. 0.001 d. 0.159 x e. 0.841
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer.

P(Z<1) = ?

Lets see the empirical value of area under normal curve.
it says that P(1<Z<1) = .68

So, P(1<Z<1) = .68
P(Z<1) = 1- (P(Z>1)+P(Z<1))/2 = .68
2*P(Z<1) = .32 [P(Z>1) = P(Z<1), since normal distribution is symmetrical]

P(Z<1) = .16

Answer is : the area under curve to the left of Z=1 is 0.16